OpenStudy (anonymous):

integrate (2x+1)/(x^2+1)

OpenStudy (anonymous):

\[\int\limits_{}^{}(2x+1)/(x ^{2}+1)dx=\int\limits_{}^{}2xdx/(x ^{2}+1)+\int\limits_{}^{}dx/x ^{2}+1\] this gives u \[\ln(x ^{2}+1)+\tan^{-1} x+c\]